OpenStudy (anonymous):

Calculate the formula mass for the following chemicals. Then solve for the appropriate number assigned. 2.45g of BrCl. Convert to mol. 4.56g of Cr(OH)3. Convert to mol .56g of NBr3. Convert to molecules. 2.2 mol of MgF2. Convert to molecules. 1.22g of lead(IV) perchlorate. Convert to molecules 5.77x 1021 molecules of dinitrogen tetroxide. Convert to grams. 7.11 x 1024molecules of manganese(IV) oxide. Convert to grams.

OpenStudy (abb0t):

I will help you with the first one. you can try the rest yourself and if you still need more help, I can guide you to get the correct answer if you're wrong. For this, you are going to use dimensional analysis. First, you look at your periodic table to get the molecular weight of bromine According to the periodic table I am looking at, it is 79.90 g, which is the same as g/ mol. Next, find the molecular weight for chlorine: 35.45 g/ mol Now, using dimensional analysis, which works similarly to how you use fractions to cross cancel out variables \[2.45 g \left( \frac{ 1 mol }{ 35.45 + 79.90 = BrCl (grams) } \right)\] notice that the 2.45 and sum of the BrCl grams cancel out. to give you mol ONLY. That is your answer of moles of BrCl.